Which one of the following will not likely occur as a result of the economy’s automatic stabilizers operating during periods of inflation?

a) The size of the investment and government spending multiplier falls.

b) The aggregate spending curve shifts upward.

c) The aggregate spending curve shifts downwards.

d) The MPC of national income is reduced.

e) The rate of inflation is slowed




Suppose the consumption function is C = 100 + 0.95YD. If the tax rate changes from t = 0 to t = 30%, then the increase in government spending that leaves the equilibrium income unaffected is:

a) 1,478

b) 2,000

c) 2,570

d) 10,280

e) 570



Other things equal, a decrease in the price level will

a) Move the economy down a given aggregate demand curve.

b) Move the economy up a given aggregate demand curve.

c) Shift the aggregate demand curve to the right.

d) Shift the aggregate demand curve to the left.

e) None of the above





1
Expert's answer
2021-12-06T09:37:22-0500

1.b) The aggregate spending curve shifts upward.


2.If t = 0, then T = 0, so C = 100 + 0.95(Y - T) = 100 + 0.95Y.

If t = 0.3, then T = t"\times" Y = 0.3Y, so C = 100 + 0.95(Y - 0.3Y) = 100 + 0.665Y

Y=570

e) 570


3.b) Move the economy up a given aggregate demand curve.